OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side in the following triangle measures. Round to the nearest tenth, if necessary. a = 5, b = ?, c = 14

OpenStudy (anonymous):

sqrt(14^2-5^2)

OpenStudy (anonymous):

13.1

OpenStudy (anonymous):

5+B=14 25+B=210 -25 -25 B=185 (Square Root) B=13

OpenStudy (anonymous):

B=13.6

OpenStudy (anonymous):

b = √( 14² - 5²) = √171 = 13.1